# TransitReID: Transit OD Data Collection with Occlusion-Resistant Dynamic Passenger Re-Identification

### v1

@@ III-B Detection and Tracking
These equations correspond to trajectories of alighting, boarding, moving while inside the bus, and moving while remaining outside. Only trajectories satisfying the first two cases are retained, with their corresponding locations recorded as key-value pairs.
@@ III-D Grading Model
Passengers’ local features should be weighted differently to account for various occlusion and viewpoint scenarios. Existing work [10] utilizes a convolutional network including a convolution layer and a fully connected layer to compute local scores. [47] designs a spatial-patch contrastive (SPC) loss to compute aggregation maps to focus on different areas. Our grading model uses a Variational Autoencoder (VAE) to assign scores to images of different body parts, where a higher score indicates better image quality, with less occlusion and a clearer viewpoint. This score will be used to calculate image weights in the subsequent processing stages.
@@
The structure of the grading model is shown in Fig. 5. The image input passes through an encoder, which consists of a convolutional layer, followed by a max-pooling layer after each convolution. The features are then processed by residual modules to obtain a flattened feature map. Two fully connected layers calculate the mean and variance of the latent space, introducing randomness to the model and enabling better learning of reconstruction features. The decoder consists of a series of transposed convolutional layers and residual modules which output the reconstructed image through a sigmoid function.
During the training, the grading model is jointly trained with the feature extraction network Global-guided Reciprocal Learning (GRL) [48]. The underlying logic is straightforward: through training, the VAE is enabled to better reconstruct images of higher quality and poorly reconstruct those with severe occlusions and bad viewpoints. The image score, $I_{S}$, is calculated based on the joint similarity between the reconstructed and original images:
$I_{S}=\sigma_{1}SSIM+\sigma_{2}HS+\sigma_{3}MSE$ (2)
To comprehensively assess the image reconstruction effect, the above formula considers three measurement factors. The Structural Similarity Index Measure (SSIM) [49] is an indicator to measure the visual similarity of two images based on three comparison parameters: luminance, contrast, and structure. Histogram Similarity (HS) focuses on the pixel value distribution of images, calculating the similarity of the histograms of the raw and reconstructed images. The Mean Squared Error (MSE) reflects the pixel value differences between the two images. The score is then fed into the Score Mapping for further process.
@@
A Logarithmic Transformation is used to map the original scores to a range from 0 to 1. Additionally, the logarithmic transformation compresses a wide score range into a smaller output range, which benefits our scenarios, where severe occlusions are of greater concern, the mapped scores decrease rapidly, helping to effectively suppress high-occlusion situations. For cases with minor occlusions, we aim to preserve the features of these images to the greatest extent. Equation (3) defines a mapping $\mathcal{L}:x\rightarrow y$, where $x$ is the input score and $y$ is the transformed score. $min/max$ is the upper/lower bound of the input. Fig. 6 displays the mapping curves at different values of $k$.
$y=\frac{\log(1+k\cdot(x-\min))}{\log(1+k\cdot(\max-\min))}$ (3)
@@ III-E Selective Quality Feature Averaging
Initially, a set of images of a passenger is extracted from a video, $\{I_{1},I_{2},...,I_{l}\}$, where $l$ denotes the length of the image sequence. These images are sequentially input into a feature extraction network, which outputs a fixed-length one-dimensional vector as the feature description of the passenger. We employ GRL[48] as the feature extraction network. It effectively utilizes both global and local feature correlations within video frames. This Global-guided Reciprocal Learning (GRL) framework incorporates two key modules: the Global-guided Correlation Estimation (GCE) and the Temporal Reciprocal Learning (TRL). The GCE module generates correlation maps to distinguish between high- and low-correlation features under global guidance, enhancing the model’s focus on relevant features. The TRL module optimizes the use of these features over time, improving the discrimination of temporal information.
The features for three body parts Head, Torso, Legs ($H,T,L$), are processed separately to obtain the feature sets $\mathfrak{F}=\{F_{H},F_{T},F_{L}\}$, where $\underset{k=H,T,L}{F_{k}}=\{f_{1}^{k},f_{2}^{k},...,f_{l}^{k}\}$. The score sets output by the grading model are denoted by $\mathcal{S}=\{S_{H},S_{T},S_{L}\}$, where $\underset{k=H,T,L}{S_{k}}=\{s_{1}^{k},s_{2}^{k},...,s_{l}^{k}\}$. We use $l=8$ frames per person for consistency.
Figure 7: Selective Quality Feature Averaging. Each image is attached with a score reflecting its degree of occlusion. Averaging all frames would suppress those of high quality. Averaging only frames with scores above a threshold, however, would preserve the features of high quality.
@@
$\mathcal{F}=\underset{k=H,T,L}{\mathtt{CONCAT}}\;(\frac{F_{k}\odot S_{k}}{l-d})$ (4)
Where $d$ is the number of images in the sequence whose scores fall below a certain threshold. $\odot$ is the Hadamard product. $\mathtt{CONCAT}$ indicates that vectors from the three parts are concatenated along the last dimension to form a one-dimensional feature vector. These features would then be stored in the Facebook AI Similarity Search (FAISS) library [50] which is an optimized library for large-scale high dimensional nearest neighbor similarity search.
@@ III-F Hierarchical Storage and Dynamic Matching
Figure 8: Structure of hierarchical storage and dynamic matching.
@@ III-F3 Final Update
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ IV Experiments
We evaluate the performance of this work from two aspects. First, we evaluate the passenger ReID on our own dataset and compare it with other state-of-the-art approaches. The training is on a single RTX4090 24GB GPU with the batch size, epochs, and initial learning rate set to 8, 60, and 0.0001 with a decay rate 0.1 after 30 epochs, respectively, while retaining some default settings for PFD [34] and TransReID [51]. Next, we run the simulation for transit OD data collection along a manually configured bus route on an NVIDIA Jetson AGX Orin. The confidence threshold of HSDM is set to 0.12 based on the confidence distribution of the training data. This value maximizes the discrimination between positive and negative samples. The logarithmic transformation parameter $k=20$. The weights $\sigma_{1},\sigma_{2},\sigma_{3}$ of the joint similarity are set to $0.2,0.6,0.2$, respectively. To speed up processing, only 8 images are randomly sampled to form a passenger’s tracklet. The loss function follows the same setting in [48]. Performance metrics include Rank-1, 5, 10, 20, mean Average Precision (mAP), and the Cumulative Matching Characteristic (CMC) curve.
IV-A Bus ReID Dataset
@@
IV-B Evaluation on ReID Performance
The performance of TransitReID on traditional ReID tasks is presented in Table I, where it is compared with several state-of-the-art methods, including PGFA [18], HOReID [52], GRL [48], TransReID [51], PFD [34], and CLIP-ReID [53]. TransitReID achieves the highest accuracy among all comparative methods, improving Rank-1 accuracy by 4.7% and mAP by 8.9%. It is noteworthy that PGFA, HOReID, TransReID, and PFD are specifically designed for occluded ReID, which explains their superior performance compared to the vision-language model CLIP-ReID. Although our baseline model GRL is not specifically designed for occluded cases, it still outperforms some occluded methods due to its video-based ReID approach that learns temporal correlations. Our method not only incorporates temporal continuity information but also selectively focuses on less occluded regions from clear viewpoints to extract useful information. The experimental results show the effectiveness of our method in handling multi-view and complex occlusion scenarios specific to passenger ReID in transit environments.
@@ IV-B Evaluation on ReID Performance
Matching Rate (%)
Rank 1 5 10 20 mAP
TransitReID (Ours) 88.3 96.7 100 100 92.5
PGFA[18] 60.9 70.2 75.3 80.4 44.3
HORelD[52] 69.0 75.6 78.2 82.0 58.8
GRL[48] 75.0 96.7 100 100 83.6
TransRelD[51] 83.6 87.1 89.1 91.4 78.8
PFD[34] 78.8 84.0 86.6 89.3 74.6
CLIP-ReID[53] 55.4 68.4 75.0 82.8 45.7
Figure 11: Cumulative Matching Characteristic curve of different ReID methods.
Fig. 11 presents the CMC curves of TransitReID and its comparative methods from Rank-1 to Rank-20. Our method achieves nearly 100% accuracy after Rank-9. Notably, the baseline model GRL attains almost identical accuracy with ours after Rank-5, demonstrating its high performance in video-based ReID tasks. However, TransitReID significantly outperforms GRL from Rank-1 to Rank-4, validating its effectiveness in suppressing occlusion interference. From another perspective, it enhances the baseline model’s matching capability on ‘hard cases’, where the target gallery person often appears within the top 5 matches but fails to achieve the best match due to difficulties in extracting key features and confusion from very similar samples.
@@
TABLE II: Performance across different numbers of bus stops without / with HSDM.
Stop ID 5 6 7 8 9 10 11 12 13 14 15
Rank-1 0.87 / 0.90 0.87 / 0.90 0.87 / 0.91 0.88 / 0.91 0.86 / 0.90 0.88 / 0.92 0.89 / 0.92 0.88 / 0.91 0.90 / 0.92 0.89 / 0.90 0.89 / 0.90
Rank-5 0.96 / 0.94 0.95 / 0.94 0.95 / 0.96 0.95 / 0.96 0.95 / 0.94 0.95 / 0.95 0.95 / 0.96 0.95 / 0.95 0.95 / 0.96 0.95 / 0.94 0.95 / 0.94
Rank-10 0.96 / 0.94 0.95 / 0.94 0.95 / 0.96 0.95 / 0.96 0.95 / 0.94 0.95 / 0.95 0.95 / 0.96 0.95 / 0.95 0.95 / 0.96 0.95 / 0.94 0.95 / 0.94
IV-C Evaluation on Dynamic Transit OD Data Collection
@@
IV-D Inference Time Analysis on Edge Devices
We conduct experiments on an NVIDIA Jetson AGX Orin to simulate the inference time of an edge computing device in a realistic transit scenario. The results are presented in Table III. Board/Alight indicates the number of passengers boarding or alighting at each stop. $T_{F}$ represents the total inference time for feature extraction, and $T_{M}$ represents the total time for ReID retrieval. The last row shows the average inference time per passenger. The processing time per passenger is approximately 3 seconds, mainly constrained by the computational performance of the edge device. Nevertheless, We implement multi-threaded programming to enable real-time operation in practical smart transit applications.
@@ IV-D Inference Time Analysis on Edge Devices
The real-time performance of the detection threads ensures smooth operation within the transit system. Meanwhile, the featurize thread does not require strict real-time execution—as long as the processing is completed before the next stop. For example, if 10 passengers board the bus at stop $i$, the entire processing should be completed within 30 seconds, before the bus arrives at stop $i+1$.
@@
TABLE III: Statistics of each stop’s information and processing time.
@@
IV-E Ablation Study
The effectiveness of the proposed HSDM mechanism is evaluated through simulations with the number of bus stops $S$ ranging from 5 to 15. Each configuration involves 20 repeated experiments, and the average matching rate across all stops is calculated. The results are shown in Table II.
Without HSDM, the Rank-1 matching rate increases very slightly as the number of bus stops increases. This is attributed to the reduction in gallery samples at each stop, which enhances the matching success rate. With HSDM, the Rank-1 rate increases across all settings, demonstrating the mechanism’s capability to handle low-confidence matches by providing these samples with additional opportunities to achieve the global optimum.
Notably, with HSDM, the Rank-1 rate remains stable regardless of the number of bus stops. This stability is achieved through the cold storage mechanism, which ensures that gallery samples at each stop are influenced not only by the number of passengers on board but also by their matching confidence. Consequently, the number of matching samples are larger than the actual number of passengers onboard, further validating the robustness of the proposed method under varying numbers of bus stops.
For Rank-5 and Rank-10, no significant changes are observed with or without HSDM, indicating the high accuracy of our ReID algorithm for simple samples.
@@ V Conclusion
This paper presented TransitReID, a novel framework for individual-level and dynamic transit OD data collection. The proposed solution addressed two critical challenges in transit environments: (1) severe occlusion through an innovative region-attention ReID mechanism that integrates a VAE-based grading model to dynamically weight visible body regions, coupled with adaptive feature fusion to preserve discriminative information; (2) continuous operation requirements via a Hierarchical Storage and Dynamic Matching (HSDM) architecture that incorporated Hot/Cold Storage and a Snatch mechanism to balance accuracy, storage, and speed. Furthermore, a multi-threaded design is implemented to achieve near real-time operation on transit vehicles equipped with edge devices. As all features are processed and stored locally, passenger privacy is effectively protected. Comprehensive evaluations demonstrated the framework’s effectiveness, achieving state-of-the-art performance with 88.3% Rank-1 accuracy on our proposed transit ReID dataset and maintaining on average 90% OD estimation accuracy in real-world transit simulations.
While real video data are now available, the configuration of bus stops remains a manual process. Future work will involve deploying this system in operational public transport vehicles to conduct further tests and refinements, ensuring real-time performance. Moreover, there is a pressing need to develop lighter algorithms based on this research to facilitate real-time execution on edge devices.

### v2

@@ III-B Detection and Tracking
These equations correspond to trajectories of 1) alighting, 2) boarding, 3) moving inside the bus, and 4) moving while remaining outside. Taking the second formula as an example, a passenger is classified as boarding only if they first appear in $R_{door}$, then leave from the $R_{inside}$ region, and do not return to $R_{door}$. This processing effectively filters out false boarding and alighting behaviors, such as when a passenger lingers near the door but ultimately does not board the bus. Only trajectories satisfying the first two cases are retained, with their corresponding locations recorded as key-value pairs.
@@ III-D Grading Model
We design a grading model using a Variational Autoencoder (VAE) to assign scores to images of different body parts, where a higher score indicates better image quality, with less occlusion and a clearer viewpoint. This score will be used to calculate image weights in the subsequent processing stages.
@@
The structure of the grading model is shown in Fig. 5. During the training, the grading model is jointly trained with the feature extraction network. The underlying logic is straightforward: through training, the VAE is enabled to better reconstruct images of higher quality and poorly reconstruct those with severe occlusions and bad viewpoints. Note that this VAE structure is inherently robust to complex occlusion patterns and dynamic viewpoints, as it relies on the holistic integrity of the image, regardless of the viewing angles or how the occlusion occurs.
The image score $I_{S}$ is calculated based on the joint similarity between the reconstructed and original images:
$I_{S}=\sigma_{1}\text{SSIM}+\sigma_{2}\text{HS}+\sigma_{3}\text{MSE}$ (6)
To comprehensively assess the image reconstruction effect, the above formula considers three measurement factors. The Structural Similarity Index Measure (SSIM) [53] measures the visual similarity of two images. The Histogram Similarity (HS) calculates the similarity of the histograms. The Mean Squared Error (MSE) reflects the pixel value differences.
@@
A logarithmic transformation is applied to map the original scores to a range of 0 to 1. It compresses a wide score range into a narrower output range, which is particularly beneficial in scenarios where severe occlusions are of greater concern. In such cases, the lower scores decrease rapidly, effectively suppressing the impact of high-occlusion situations. Conversely, for images with minor occlusions, the transformation preserves their features to the greatest extent. Equation (3) defines a mapping $\mathcal{L}:x\mapsto y$, where $x$ is the score and $\min/\max$ is the upper/lower bound of the input. Fig. 6 displays the mapping curves at different values of $k$.
$\mathcal{L}(x)=\frac{\log(1+k\cdot(x-\min))}{\log(1+k\cdot(\max-\min))}$ (7)
@@ III-E Selective Quality Feature Averaging
Initially, a set of images of a passenger is extracted from a video, $\{I_{1},I_{2},...,I_{l}\}$, where $l$ denotes the length of the image sequence. These images are sequentially input into a feature extraction network, which outputs a fixed-length one-dimensional vector as the feature description of the passenger. We employ GRL[54] as the feature extraction network. It effectively utilizes both global and local feature correlations within video frames.
The features for three body parts Head, Torso, Legs ($H,T,L$), are processed separately to obtain the feature sets:
$\mathcal{F}=\{F_{k}=\{f_{1}^{k},f_{2}^{k},\dots,f_{l}^{k}\}\}_{k\in\{H,T,L\}}$ (8)
The score sets output by the grading model are denoted by:
$\mathcal{S}=\{S_{k}=\{s_{1}^{k},s_{2}^{k},\dots,s_{l}^{k}\}\}_{k\in\{H,T,L\}}$ (9)
We use $l=8$ frames per person for consistency.
Figure 7: Selective quality feature averaging (SQFA). Each image is attached with a score reflecting its degree of occlusion. Averaging only frames with scores above the threshold (as shown in the boxes) would preserve the features of high quality.
@@
$\widetilde{F}=\operatorname{CONCAT}_{k\in\{H,T,L\}}\Biggl{(}\frac{F_{k}\odot S_{k}}{l-d}\Biggr{)}$ (10)
Where $d$ is the number of images in the sequence whose scores fall below a certain threshold. $\odot$ is the Hadamard product. $\operatorname{CONCAT}$ indicates that vectors from the three parts are concatenated along the last dimension to form a one-dimensional feature vector. These features would then be stored in the Facebook AI Similarity Search (FAISS) library [55] which is an optimized library for large-scale high dimensional nearest neighbor similarity search.
@@ III-F Hierarchical Storage and Dynamic Matching
Figure 8: Structure of hierarchical storage and dynamic matching (HSDM).
@@ III-F3 Final Update
6:   $(\text{topK\_IDs},\text{topK\_dists})=\text{FAISS.search}$
@@
10:     $p_{\text{alight}}.\text{matched\_ID}=\text{topK\_IDs}[0]$
11:     Remove $\text{topK\_IDs}[0]$ from FAISS index
@@
14:     $p_{\text{alight}}.\text{temp\_match}=\text{topK\_IDs}[0]$
@@
18:  $(\text{topK\_IDs2},\text{topK\_dists2})=\text{FAISS.search}$
@@
20:  if $\text{topK\_IDs2}[0]$ is in Cold Storage by another $p_{\text{alight}}$ then
@@
22:     Remove $\text{topK\_IDs2}[0]$ from $p_{\text{alight}}.\text{temp\_match}$
@@
24:      $p_{\text{alight2}}.\text{matched\_ID}=\text{topK\_IDs2}[0]$
25:      Remove $\text{topK\_IDs2}[0]$ from FAISS index
@@
27:      $p_{\text{alight2}}.\text{temp\_match}=\text{topK\_IDs2}[0]$
@@
34:  $p_{\text{alight}}.\text{matched\_ID}\leftarrow p_{\text{alight}}.\text{temp\_match}$
@@ IV Experiments
IV-A Implementations
We evaluate the performance of this work from two aspects. First, we evaluate the passenger ReID on our own dataset and compare it with other state-of-the-art approaches. The training is on a single RTX4090 24GB GPU with the batch size, epochs, and initial learning rate set to 8, 60, and 0.0001 with a decay rate 0.1 after 30 epochs, respectively, while retaining some default settings for PFD [40] and TransReID [56]. Next, we run the simulation for transit OD data collection along a manually configured bus route on an NVIDIA Jetson AGX Orin. The confidence threshold of HSDM is set to 0.12 based on the confidence distribution of the training data. This value maximizes the discrimination between positive and negative samples. The logarithmic transformation parameter $k=20$. The weights $\sigma_{1},\sigma_{2},\sigma_{3}$ of the joint similarity are set to $0.2,0.6,0.2$, respectively. To speed up processing, only 8 images are randomly sampled to form a passenger’s tracklet. The loss function follows the same setting in [54]. Performance metrics include R-1, 5, 10, 20 and mean Average Precision (mAP)
IV-B Bus ReID Dataset
@@
IV-C Evaluation on ReID Performance
The performance of TransitReID on traditional ReID tasks is presented in Table I, where it is compared with several state-of-the-art methods, including PGFA [17], HOReID [57], GRL [54], TransReID [56], PFD [40], CLIP-ReID [58], and TF-CLIP [59]. TransitReID achieves the highest accuracy among all comparative methods, improving R-1 accuracy by 1.6% and mAP by 1.8%. It is noteworthy that PGFA, HOReID, and PFD are specifically designed for occluded ReID, yet our method still achieves better results. Although our baseline model GRL is not specifically designed for occluded cases, it still outperforms some occluded methods due to its video-based approach that learns temporal correlations. Our method not only incorporates temporal information but also selectively focuses on less occluded regions from clear viewpoints to extract fine-grained information. The experimental results show the effectiveness of our method in handling multi-view and complex occlusion scenarios specific to passenger ReID in transit environments.
@@ IV-C Evaluation on ReID Performance
Methods References R-1 R-5 R-10 R-20 mAP
PGFA[17] ICCV’2019 60.9 70.2 75.3 80.4 44.3
HORelD[57] CVPR’2020 69.0 75.6 78.2 82.0 58.8
GRL[54] CVPR’2021 75.0 96.7 100 100 83.6
TransRelD[56] ICCV’2021 83.6 87.1 89.1 91.4 78.8
PFD[40] AAAI’2022 78.8 84.0 86.6 89.3 74.6
CLIP-ReID[58] AAAI’2023 55.4 68.4 75.0 82.8 45.7
TF-CLIP[59] AAAI’2024 86.7 96.7 98.3 100 90.7
TransitReID Ours 88.3 96.7 100 100 92.5
Notably, GRL attains almost identical accuracy with ours after R-5, demonstrating its strong performance in video-based ReID tasks. However, TransitReID significantly outperforms GRL in R-1 and mAP, validating its effectiveness in suppressing occlusion interference. From another perspective, it enhances the baseline model’s matching capability on ‘hard cases’, where the target gallery person often appears within the top 5 matches but fails to achieve the best match due to difficulties in extracting key features and confusion from very similar samples.
@@
IV-D Evaluation on Dynamic Transit OD Data Collection
@@
IV-E Inference Time Analysis on Edge Devices
We conduct experiments on an NVIDIA Jetson AGX Orin to simulate the inference time of an edge computing device in a realistic transit scenario. The results are presented in Table II. Board/Alight indicates the number of passengers boarding or alighting at each stop. $T_{F}$ represents the total inference time for feature extraction, and $T_{M}$ represents the total time for ReID retrieval. The last row shows the average inference time per passenger. The processing time per passenger is about 3 seconds, mainly constrained by the computational performance of the edge device. Nevertheless, We implement multi-threaded programming to enable real-time operation in practical smart transit applications.
@@ IV-E Inference Time Analysis on Edge Devices
The real-time performance of the detection threads ensures smooth operation within the transit system. Meanwhile, the featurize thread does not require strict real-time execution as long as the processing is completed before the next stop. For example, if 10 passengers board the bus at stop $i$, the entire processing should be completed within 30 seconds, before the bus arrives at stop $i+1$.
@@
TABLE II: Statistics of each stop’s information and processing time.
@@
TABLE III: Performance across different numbers of bus stops without / with HSDM.
Stop ID 5 6 7 8 9 10 11 12 13 14 15
R-1 0.87 / 0.90 0.87 / 0.90 0.87 / 0.91 0.88 / 0.91 0.86 / 0.90 0.88 / 0.92 0.89 / 0.92 0.88 / 0.91 0.90 / 0.92 0.89 / 0.90 0.89 / 0.90
R-5 0.96 / 0.94 0.95 / 0.94 0.95 / 0.96 0.95 / 0.96 0.95 / 0.94 0.95 / 0.95 0.95 / 0.96 0.95 / 0.95 0.95 / 0.96 0.95 / 0.94 0.95 / 0.94
R-10 0.96 / 0.94 0.95 / 0.94 0.95 / 0.96 0.95 / 0.96 0.95 / 0.94 0.95 / 0.95 0.95 / 0.96 0.95 / 0.95 0.95 / 0.96 0.95 / 0.94 0.95 / 0.94
IV-F Ablation Study
HSDM
The effectiveness of the proposed HSDM mechanism is evaluated through simulations with the number of bus stops ranging from 5 to 15. Each configuration involves 20 repeated experiments, and the average matching rate across all stops is calculated. The results are shown in Table III.
Without HSDM, the R-1 matching rate increases very slightly as the number of bus stops increases. This is attributed to the reduction in gallery samples at each stop, which enhances the matching success rate. With HSDM, the R-1 rate increases across all settings, demonstrating the mechanism’s capability to handle low-confidence matches by providing these samples with additional opportunities to achieve the global optimum.
Notably, with HSDM, the R-1 rate remains stable regardless of the number of bus stops. This stability is achieved through the cold storage mechanism, which ensures that gallery samples at each stop are influenced not only by the number of passengers on board but also by their matching confidence. Consequently, the gallery size is larger than the actual number of passengers onboard, further validating the robustness of the proposed method under varying numbers of bus stops.
For R-5 and R-10, no significant changes are observed with or without HSDM, indicating the high accuracy of our ReID algorithm for simple samples.
Contributions from Different Components
We evaluate the contribution of different components on the proposed bus ReID dataset, as shown in Table IV. First, we remove one of the three body parts individually, and observe that both mAP and R-1 decrease compared to using the full body. However, when the Legs part is missing, R-5 improves, which may be because some hard cases involving legs tend to cause confusion. “G” denotes the grading model, which brings improvements of 5.2% in mAP and 10% in R-1. SQFA slightly improves R-1 by 0.1% and R-5 by 1.7%.
TABLE IV: Ablation study on the effectiveness of each component on Bus ReID dataset.
Head Legs Torso G SQFA mAP R-1 R-5
84.8 75.0 96.7
88.1 80.0 98.3
90.0 85.0 95.0
89.9 83.3 96.7
87.2 78.3 95.0
92.4 88.3 95.0
92.5 88.3 96.7
@@ V Conclusion
This paper presented TransitReID, a novel framework for individual-level and dynamic transit OD data collection. The proposed solution addressed two critical challenges in transit environments: (1) severe occlusion through an innovative region-attention ReID mechanism that integrates a VAE-based grading model to dynamically weight visible body regions, coupled with adaptive feature fusion to preserve discriminative information; (2) continuous operation requirements via a Hierarchical Storage and Dynamic Matching (HSDM) architecture that incorporated Hot/Cold Storage and a Snatch mechanism to balance accuracy, storage, and speed. Furthermore, a multi-threaded design is implemented to achieve near real-time operation on transit vehicles equipped with edge devices. As all features are processed and stored locally, passenger privacy is effectively protected. Comprehensive evaluations demonstrated the framework’s effectiveness, achieving state-of-the-art performance with 88.3% R-1 accuracy on our proposed transit ReID dataset and maintaining on average 90% OD estimation accuracy in real-world transit simulations.
While real video data are now available, the configuration of bus stops remains a manual process. Future work will involve deploying this system in public transport vehicles to conduct further tests and refinements, ensuring real-time performance. Moreover, there is a pressing need to develop lighter algorithms based on this research to facilitate real-time execution on edge devices.